### v1

# README: Bridging Medical Jargon and Lay Understanding for Patient Education through Data-Centric NLP

Abstract
The advancement in healthcare has shifted focus toward patient-centric approaches, particularly in self-care and patient education, facilitated by access to Electronic Health Records (EHR). However, medical jargon in EHRs poses significant challenges in patient comprehension. To address this, we introduce a new task of automatically generating lay definitions, aiming to simplify complex medical terms into patient-friendly lay language. We first created the README dataset, an extensive collection of over 20,000 unique medical terms and 300,000 mentions, each offering context-aware lay definitions manually annotated by domain experts. We have also engineered a data-centric Human-AI pipeline that synergizes data filtering, augmentation, and selection to improve data quality. We then used README as the training data for models and leveraged a Retrieval-Augmented Generation (RAG) method to reduce hallucinations and improve the quality of model outputs. Our extensive automatic and human evaluations demonstrate that open-source mobile-friendly models, when fine-tuned with high-quality data, are capable of matching or even surpassing the performance of state-of-the-art closed-source large language models like ChatGPT. This research represents a significant stride in closing the knowledge gap in patient education and advancing patient-centric healthcare solutions  [^1] Our code and part of data will be released at https://github.com/seasonyao/NoteAid-README.
1 Introduction
The advancement in healthcare transcends medical breakthroughs, encompassing an increased emphasis on patient involvement in self-care Boling (2009); Spooner et al. (2019); Baldry et al. (1986); Schillinger et al. (2009). The criticality of effective patient education is underscored in an era marked by the emergence of new treatments that demand patient comprehension and cooperation. Access to Electronic Health Records (EHR) has become an integral part of this educational paradigm, with initiatives like PCASSO Masys et al. (2002) and NoteAid Chen et al. (2018); Kwon et al. (2022) leading the way in empowering patients and enhancing healthcare outcomes. These efforts demonstrate a shift towards patient-centric healthcare, where informed patients play an active role in their treatment processes.
Despite advancements in electronic medical record management, one significant barrier persists in the form of medical jargon in EHRs, impeding patient understanding and self-care Kwon et al. (2022). As shown in Figure 1, tools like NoteAid, which employs Natural Language Processing (NLP) to demystify complex medical terms, have been instrumental in bridging the communication gap between healthcare professionals and patients Lalor et al. (2018, 2021, 2023). However, existing resources like the Consumer Health Vocabulary (CHV) Zeng and Tse (2006); He et al. (2017); Ibrahim et al. (2020) are limited in scale, posing a challenge to NoteAid. For instance, only a fraction (about 4%) of the medical terms in NoteAid have been annotated with lay definitions, highlighting the need for a more scalable solution to address this knowledge gap effectively.
Addressing this issue requires shifting our focus to online health education materials such as Unified Medical Language System (UMLS)  Bodenreider (2004), MedlinePlus Patrias and Wendling (2007), Wikipedia, and Google. However, as indicated in Figure 1, these resources often present information that is too difficult for the average patient to understand. For example, these resources’ average readability measured by the Flesch Kincaid Grade Level was post-secondary or higher education, while the average readability of a US adult was 7-8th grade level Doak et al. (1996, 1998); Eltorai et al. (2014). To bridge this gap, we have engaged a team of medical experts to meticulously curate lay definitions for jargon terms found in NoteAid, targeting a comprehension level suitable for individuals with a 7th to 8th-grade education. Each term within the NoteAid dataset has been redefined across various contexts, ensuring their applicability in diverse clinical scenarios. This effort led to the creation of the REsource of lAy Definitions for MEdical jargon (README) dataset, an expansive resource containing lay definitions for over 20,000 medical terms. In total, the README dataset comprises an impressive 300,000 data points, each consisting of a clinical note context, a medical jargon term, and its corresponding lay definition, thereby significantly enhancing the accessibility and comprehensibility of medical information for patients.
Figure 1: The NoteAid pipeline comprises an NLP component and a resource of medical jargon and the corresponding lay definitions. We list definitions for the term "EGD" across various online sources registered Flesch-Kincaid Grade Level (FKGL) above 13, suggesting they are comprehensible mainly to individuals with post-secondary education.
Figure 2: Our Human-AI-in-the-loop Data-Centric NLP pipeline, comprising the Examiner-Augmenter-Examiner (EAE) framework and different data selection methods. EAE shows how humans (physicians) and AI (LLM, e.g. ChatGPT) cooperate to make a high-quality README dataset. We collect general definitions for every jargon term from external knowledge resources such as UMLS. “R” is “README”. “v1” is “version-1”. “instruction” and “demo” (examples for in-context learning) are combined into the prompt for LLM. In the pipeline, the human duties at different stages are annotator (labeling the initial dataset) and instructor (providing suitable prompts to guide AI at every stage). The AI duties at different stages are examiner (filter high-quality data) and augmenter (improve the quality of low-quality data). We then deploy 4 different data selection strategies to combine high-quality expert-annotated data R-v2 and high-quality AI-synthetic data R-v5 and train the in-house system.
Yet, the critical aspect of generating lay definitions remains largely unexplored. As patients gain more access to their EHRs, the demand for lay definition resources is escalating, inevitably destined to surpass the capacity of current expert-annotated resources, regardless of efforts to expand them. In addition, the dynamic nature of "jargon" based on individual and context makes pre-annotated expert resources less adaptable to real-life scenarios. The model-driven automatic generation of lay definitions from medical jargon emerges as a viable solution. Recent research highlighted ChatGPT’s potential in its integration with the field of medicine Brown et al. (2020); OpenAI (2023); Yang et al. (2023), including generating human-readable definitions for biomedical terms Remy and Demeester (2023). Nonetheless, our evaluations of open-source models (refer to Figure 4) indicate notable differences in their performance compared to ChatGPT, particularly in terms of medical knowledge Sung et al. (2021); Yao et al. (2022, 2023).
To bridge this gap, we aim to train an in-house system using open-source models for automatic lay definition generation to provide reliable lay definitions for jargon in patient education tools like NoteAid. Inspired by research on Retrieval-Augmented Generation (RAG)  Lewis et al. (2020); Asai et al. (2023), we aim to overcome the limitations of open-source base models in medical knowledge. We are positioning automatic lay definition generation as a form of text simplification, where language models are prompted to generate context-aware, jargon-specific, and layperson-friendly definitions based on standard definitions retrieved from external knowledge resources. Specifically, in this work, we use the UMLS to retrieve standard definitions of jargon terms and construct a dataset upon the README that includes context, jargon terms, standard definitions, and lay definitions.
Subsequently, we designed the data-centric pipeline Examiner-Augmenter-Examiner (EAE) to enhance the data quality within the README dataset, as seen in Figure 2. Specifically, following a human-in-the-loop paradigm Monarch (2021), we employed human experts to guide AI in both examiner and augmenter stages, where the former selects high-quality training data (originating from either manual annotation or AI synthesis), and the latter generates potentially high-quality synthesized data to increase data points. Finally, we employed the AI-synthesized dataset to augment the expert-annotated dataset, aiming to explore the effectiveness of AI-generated data in training, especially in scenarios with limited expert data. We implemented a range of heuristic data selection strategies to integrate AI synthetic data, allowing us to incorporate suitable data points into our training process.
In summary, our contributions are as follows:
  • [leftmargin=.2in,topsep=0.1pt]
  • Introduced a new task of automatically generating lay definitions for medical jargon. We created a substantial expert-annotated dataset containing 300,000 data points, serving as a detailed dictionary and a benchmark for this novel task. We leveraged a Retrieval-Augmented Generation (RAG) method, which aims to reduce hallucinations and improve the quality of model outputs. Our work aims to enhance patient education by improving patients’ comprehension of medical documentation.
  • Developed a robust, data-centric pipeline that effectively integrates data filtering, augmentation, and the selection of synthetic data. This approach enhances the quality of README datasets, merging the strengths of AI with human expertise to achieve optimal results.
  • Our extensive automatic and human evaluations reveal that when trained with high-quality data, open-source, mobile-friendly small models can achieve or even exceed the performance of cutting-edge closed-source large language models, such as ChatGPT.
2 Related Work
Data-Centric AI In the Data-centric AI framework, the README meticulously navigates through the collection, labeling, preparation, reduction, and augmentation phases Zha et al. (2023); Ng et al. (2021). README begins with data collection and labeling (4.1), harnessing domain expertise to curate a dataset that is both expansive and representative Whang et al. (2023). During the preparation phase (4.2), raw data undergoes rigorous cleaning and transformation, readying it for effective model training Krishnan and Wu (2019). README also adeptly applies data augmentation techniques (4.2) to enrich the dataset with verified quality AI-synthetic data Chlap et al. (2021). Finally, through data reduction (4.3), README selects more suitable instances from AI-synthetic data for data integration with expert-annotated data Whang et al. (2023). Collectively, these stages underscore README’s comprehensive application of Data-centric AI in the healthcare domain.
Patient Education Patient education plays a crucial role in the success of therapy (Bastable, 2016; Golper, 2001). Elevating the level of patient education has consistently been an important component within the healthcare system (McCarthy et al., 2013). By providing clear and easily understandable medical information, patient education not only enhances patients’ awareness of their own health conditions but also encourages them to actively participate in medical decision-making and self-management(Gruman et al., 2010; Coulter, 2012). NoteAid (Chen et al., 2018) transforms complex medical terms in EHR notes into understandable language, integrating advanced features like MedJEx (Kwon et al., 2022) to enhance patient comprehension. Efforts to demystify clinical language have also been made in the works of  Petrova (2014); Petrova et al. (2015), while Remy and Demeester (2023) successfully employed ChatGPT for generating accurate medical definitions.
3 Problem Statement
Consider a dataset $D=\{X,Y,Z_{+}\}$ comprising $t$ EHRs, where $X=\{x^{1},x^{2},\ldots,x^{t}\}$ represents the contexts of these EHRs, $Y=\{y^{1},y^{2},\ldots,y^{t}\}$ denotes the corresponding jargon terms, and $Z_{+}=\{z_{+}^{1},z_{+}^{2},\ldots,z_{+}^{t}\}$ are the ground truth expert lay definitions. Each EHR context $x^{i}$ is a sequence of $n$ tokens, expressed as $x^{i}=\{x_{1}^{i},x_{2}^{i},\ldots,x_{n}^{i}\}$, and each lay definition $z_{+}^{i}$ consists of $m$ tokens, given by $z_{+}^{i}=\{z_{+,1}^{i},z_{+,2}^{i},\ldots,z_{+,m}^{i}\}$. The README lay definition generation task $T$ aims to train a reference model $M_{ref}$ such that $M_{ref}(z_{+}^{i}\mid x^{i},y^{i})$ is optimized. The standard approach for fine-tuning $M_{ref}$ on $T$ involves using the cross-entropy loss $\ell_{ce}(z_{+}^{i},M_{ref}(x^{i},y^{i}))$ over the dataset $D$. To enhance the training of $M_{ref}$, we introduce an additional set of general definitions $Z_{-}=\{z_{-}^{1},z_{-}^{2},\ldots,z_{-}^{t}\}$, where each $z_{-}^{i}$ corresponds to the general definition of the jargon term $y^{i}$, generated using openly available data sources (UMLS) or GPT-3.5-turbo. Our proposed EAE pipeline is designed to acquire high-quality general definition data $Z_{-}$, culminating in the augmented dataset $D_{simp}=\{X,Y,Z_{+},Z_{-}\}$. The README lay definition generation task T is then formalized as a text simplification task, where $M_{ref}$ is trained to produce $Z_{+}$ based on $X,Y,$ and $Z_{-}$. This process utilizes a selected subset $D_{SEL}\subseteq D_{simp}$, chosen according to one of the selection criteria: RANDOM, SYNTAX, SEMANTIC, or MODEL.
4 Method
4.1 Dataset Construction
In this study, we utilized a collection of EHRs sourced from an anonymized institution. The identification of medical jargon within these EHRs was performed using the MedJEx, as detailed in Kwon et al. (2022). Domain experts provided lay definitions for these jargon terms, forming the dataset known as README. For training our models, it was necessary to define these jargon terms in a general yet scientifically accurate manner. To this end, we employed UMLS to generate general definitions. UMLS is a set of files and software that brings together many health and biomedical vocabularies and standards to enable interoperability between computer systems. We used the Scispacy library to access UMLS definitions of jargon terms. we used the en_core_sci_lg model to obtain the data. we tried other models in Scispacy but found all the other models to produce the same result as the en_core_sci_lg model  [^2] https://github.com/allenai/scispacy.
Given the complexity of medical terminology, it’s noteworthy that some terms in UMLS are associated with multiple definitions. This reflects the reality that the meaning and application of a medical term can vary based on its contextual use. To accurately select the most appropriate general definition for each context, we utilized the SentenceBERT similarity score, as proposed by Reimers and Gurevych (2019), to identify the most fitting definition.
Jargon terms in our dataset varied in length and composition, with not all words necessarily having corresponding UMLS definitions. This variability necessitated the development of distinct approaches for different scenarios:
  1. [topsep=0.5pt,itemsep=0.2ex,partopsep=0.2ex,parsep=.20ex]
  2. 1.
    Single-word Terms: In cases where the jargon term comprised a single word, we either found a UMLS definition or none. Data points lacking a UMLS definition in this category were excluded.
  3. 2.
    Two-word Terms: For jargon terms composed of two words (word1 and word2), we considered several subcases:
    1. (a)
      Only word1 has a UMLS definition.
    2. (b)
      Only word2 has a UMLS definition.
    3. (c)
      Both word1 and word2 have UMLS definitions.
    4. (d)
      The phrase “word1 word2” has a collective UMLS definition.
  4. 3.
    Terms with More than Two Words: Similar to the case with two words, we can have cases where few words have UMLS definitions in the jargon term and few do not.
Figure 3: General definitions of two jargon examples.
Our solution to address these scenarios involved a unified approach, given the impracticality of tackling each case individually. As illustrated in Figure 3, we extracted the UMLS definitions for all phrases within a jargon term. This process yielded a list of strings, providing a comprehensive general definition for each term. There might not be a UMLS definition for all the words in the jargon term. In that case, we are concatenating the word directly. Subsequently, we applied LLMs for the necessary data-cleaning steps.
The initial README dataset has a lot of unusable data. There are many reasons why a data point might not be useful. The main reasons are (i) the lay definition provided by the annotator is not correct or missing. (ii) The jargon term is missing. (iii)The general definition obtained from UMLS is too scientific or not relevant to the jargon term. Using this data for training would be bad for the training as these data points will make the model worse. So these data points are not considered. This preliminary cleaning brings out data from 350K to 308K.
4.2 Examiner-Augmenter-Examiner (EAE)
4.2.1 Examiner (expert-annotated data)
Initially, basic data cleaning, as outlined in Section 4.1, was applied. To enhance this, we employed GPT-3.5-turbo, using a few-shot learning approach with seven examples - four demonstrating acceptable data points and three showing unacceptable ones. These prompts served as the ‘Human’ element in our Human-AI-in-the-loop model, as depicted in Figure 2 and detailed in Appendix Algorithm 1. The prompts are detailed in Appendix Table 5. We choose GPT-3.5-turbo here because recent work Remy and Demeester (2023) and our evaluation (4.2.4 and 5.3) show that the definitions it generates for medical terms can reach a human-satisfying level. Post-cleaning, approximately 40% of UMLS general definitions were deemed suitable for model training. The suitable data points were archived in README-v2, while the unsuitable ones were stored in README-v3. To minimize computing costs, we excluded EHRs and took all the unique data points. This was done because, in many cases, we have the same jargon term, general definition from UMLS, and lay definition but different EHR’s. This reduced our dataset from 308,239 to 51,623 data points. Upon processing these through GPT-3.5-turbo, 11,765 were classified as ‘good’ quality general definitions, and 39,856 as ‘bad’. A subsequent SQL join with contextual data removed earlier resulted in 113,659 ‘good’ and 194,580 ‘bad’ general definitions. After eliminating duplicates, the final count for bad general definitions in README-v3 was 177,140.
README-EAE
       inputs : README-v1
       output : Final README-v6 dataset with best general definitions and lay definitions
       $Initialize:$ $exam_{p}=examinerpromptforgpt3$;
       $aug_{p}=augmenterpromptsforChatGPT$;
       foreach $datapoint:README-v1$ do
            if $exam_{p}(datapoint)==yes$ then
                   README-v2.add(datapoint);
             end if
            else
                   README-v3.add(datapoint);
             end if
       end foreach
      foreach $datapoint:README-v3$ do
            $temp=aug_{p}(datapoint)$;
             README-v4.add(temp);
             if $exam_{p}(temp)==yes$ then
                   README-v5.add(temp);
             end if
            else
                   README-v6.add(temp);
             end if
       end foreach
      return README-v6;
Algorithm 1 Algorithm for Data Cleaning
4.2.2 Augmenter
Given the low yield of usable UMLS definitions, we employed ChatGPT, a recent LLM by OpenAI, to augment our dataset. The augmentation process, a critical part of our Data-centric pipeline, involved the system prompt: "Generate a general definition of the term for a professional medical student." Accompanied by two examples(Table 5), this step aimed to create complex yet general definitions suitable for professional medical students. The outcome of this process was README-v4, containing 171,831 newly generated definitions.
4.2.3 Examiner (AI-synthetic data)
The ChatGPT-generated definitions underwent a second cleaning round using the same methodology as in Examiner (expert-annotated data). Here, approximately 56% of the ChatGPT definitions were found suitable for model training, with the remaining being either contextually inappropriate or incompatible with the expert-provided lay definitions. The final tally was 96,585 ‘good’ and 75,068 ‘bad’ general definitions, stored in README-v5 and README-v6, respectively.
4.2.4 Data Quality
The final stage involved preparing the training data, which comprised good definitions from README-v2 and README-v5. This data was split into three categories: human examination data (also be used as final test data since medical experts examine this split), out-of-distribution (OOD) evaluation data, and training data, where we make sure the medical jargon in human examination and OOD evaluation will not appear in the training split. The human examination split consisted of 1000 unique terms, each accompanied by general definitions to be rated based on two criteria:
  1. [topsep=0.5pt,itemsep=0.2ex,partopsep=0.2ex,parsep=.20ex]
  2. 1.
    Hard Correlation: Marked ‘Yes’ if the lay definition closely rephrases or shares significant wording with the general definition, implying comprehensibility without advanced medical knowledge.
  3. 2.
    Soft Correlation: Marked ‘Yes’ if the general definition accurately represents the term but is slightly contextually misaligned; marked ‘No’ if the definition is incorrect or overly verbose, complicating the derivation of a lay definition.
Here is one example for Term ‘von Willebrand disease’:
Expert Definition(lay definition): A bleeding disorder. It affects the blood’s ability to clot
General definition: Hereditary or acquired coagulation disorder characterized by a qualitative or quantitative deficiency of the von Willebrand factor. The latter plays an important role in platelet adhesion. Signs and symptoms include bruises, nose bleeding, gum bleeding following a dental procedure, heavy menstrual bleeding, and gastrointestinal bleeding.
Although the lay definition is not incorrect, it is very wordy and complex and matches less to the lay definition. We consider this soft correlated and not hard correlated.
A Hard Correlation automatically implies a Soft Correlation. Our findings revealed that 88% of README-v2 and 100% of README-v5 met the Hard Correlation criteria. There is a 100% soft Correlation criteria for both README-v2 and README-v5.
4.3 Integration of Synthetic and Expert Data
We adopted four distinct sampling strategies to integrate the AI-synthetic README-v5 data with the expert-annotated README-v2 dataset:
  • [leftmargin=.2in,topsep=2pt]
  • RANDOM: This approach randomly selected N entries from the README-v5 dataset. This is the baseline for our subsequent three heuristic methods.
  • SYNTAX: For the syntax-based sampling approach, the ROUGE_L metrics F1 score in Section 5.1 was used as a key evaluative tool. ROUGE_L focuses on the longest common subsequence, which measures the longest string of words that occurs in both the predicted and reference texts. By using this metric, we could rank the synthetic definitions according to their syntactic closeness to the human-written definitions, which helped us select samples that would potentially be more understandable and natural-sounding.
  • SEMANTIC: For semantic-based sampling, we utilized the all-MiniLM-L6-v2 model from the SentenceTransformers framework  [^3] https://github.com/UKPLab/sentence-transformers. Renowned for its text semantic analysis efficiency, this model enabled us to measure the semantic similarity between lay definitions in README-v2 and README-v5 datasets. We ranked synthetic data based on these scores, considering higher scores as indicative of greater semantic closeness to expert annotations.
  • MODEL: In model-based sampling, we used models initially trained on the README-v2 dataset to generate definitions for the README-v5 dataset. This tested the model’s capacity to transfer learnings from expert data to synthetic data. We employed the ROUGE_L F1 score to evaluate the alignment between model-generated and actual README-v5 lay definitions. This technique aids in mitigating training challenges associated with data heterogeneity and enriches the dataset with examples that enhance the model’s convergence towards the desired distribution (e.g., expert-annotated lay definitions).
5 Experiments
5.1 Automatic Evaluation Metrics
Models are evaluated with full-length F1-scores of ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005). We use QuickUMLS [^4] https://github.com/Georgetown-IR-Lab/QuickUMLS to extract medical concepts from both model-generated and ground truth summaries and then calculate F1-scores for these two lists of concepts, which is named UMLS-F1 Adams et al. (2023).
5.2 Experimental Setting
We use the following symbols:
  1. 1.
    jargon2lay(J2L): Directly generates a lay definition for a given jargon term.
  2. 2.
    jargon+context2lay(J+C2L): Generates a lay definition for a given jargon term based on the context information from clinical documents.
  3. 3.
    jargon+gen2lay(J+G2L): Generates a lay definition for a given jargon term based on the general definition provided by UMLS.
  4. 4.
    jargon+context+gen2lay(J+C+G2L): Generates a lay definition for a given jargon term based on both the context information from clinical documents and the general definition provided by UMLS.
We use the following base models: GPT-2  [^5] https://huggingface.co/gpt2 Radford et al. (2019), DistilGPT2  [^6] https://huggingface.co/distilgpt2, BioGPT  [^7] https://huggingface.co/microsoft/biogpt Luo et al. (2022), and Llama2  [^8] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023) in our experiments. We trained the base models on the different README dataset variants with Supervised Fine-tuning for 100000 steps (batch size 8)  [^9] We did all the experiments with 1 NVIDIA Tesla RTX 8000 GPU - 40 GB memory, with Adam optimizer – betas=(0.9,0.999), epsilon=1e-08, learning rate=5e-04.. In all our evaluations, we used a beam size of 4, no-repeat-ngram-size=2, and minimum length and maximum length of sentences were set as (10, 100). We used five different random seeds to sample training data for all our experiments, and the scores reported in the tables are the average of these random seeds.
In this task, we ask for your expertise in generating the corresponding lay definition from the medical jargon. Mainly, we provide the target medical jargon term. We need you to generate a lay definition for this jargon term.
Example:
jargon term: [TERM]
lay definition: [DEFINITION]
jargon term: [TERM]
lay definition:
Table 1: One shot prompt for experiment set-1.
Our experiments were divided into five distinct parts. In Set-1, we aimed to evaluate the performance gap on the jargon2lay task between open-source base models and GPT models. We do one-shot prompting for Llama2, GPT-3.5-turbo, and GPT-4 with prompts in Table 1. Set-2 explored the varying data quality across different versions within our EAE pipeline, where we fine-tuned the GPT-2 model on the jargon+gen2lay task. Set-3 focused on evaluating the effects of various data selection strategies on data augmentation outcomes. To do this, we ranked AI-generated data (v5) using different methods in Section 4.3, selecting either the top N entries with the highest scores (e.g., ‘SEMANTIC’ in Table 3) or the bottom N entries with the lowest scores (e.g., ‘SEMANTIC_r’). These selections were then evenly combined with expert-annotated data (v2) at a one-to-one ratio. Following this, we fine-tuned the GPT-2 model using these diverse, mixed datasets to determine the impact of each selection method on model performance. In Set-4, we investigated the effects of incorporating different types of information (medical documentation context and UMLS-retrieved general definition) into the model inputs. In Set-5, we fine-tuned models of varying sizes (ranging from DistilGPT2-88M to Llama2-7B) and compared the outcomes with GPT-3.5-turbo. The next section (Section 6) further validates the findings obtained through this section’s automatic evaluation with a comprehensive human evaluation approach.
5.3 Results
We start with our experiments with some base models’ performance on the README jargon2lay dataset with In Set-1. In Section 4.2.4, we noted a high level of agreement between human evaluators and definitions generated by GPT-3.5-turbo. This aligns with the findings of recent work by Remy and Demeester (2023), which underlines GPT-3.5-turbo’s efficacy in crafting human-readable explanations for biomedical concepts. Consequently, GPT-3.5-turbo serves as a strong baseline for quality in README lay definition generation, indicative of a standard that meets human satisfaction. Despite this, our assessments of Llama2, depicted in Figure 4, show a notable gap in performance. This disparity underscores the urgent need to upgrade the performance of Llama2 and other open-source models (e.g., DistilGPT2, GPT2, BioGPT) to align with the high-quality output of GPT-3.5-turbo. Finally, We also observed that GPT-3.5-turbo and GPT4 performed equally well in this task. Next, we discuss how different modules of our data-centric pipeline improve the performance of the fine-tuned models in several subsections.
Figure 4: One-shot performances on jargon2lay.
5.3.1 Effectiveness of EAE pipeline
J+G2L ROUGE1 ROUGE2 ROUGEL METEOR UMLS-F1 Rank
v1 23.94 8.95 22.79 17.75 12.83 2
v2 26.99 10.76 25.57 21.19 16.88 4
v3 22.25 7.46 21.09 17.18 11.72 5
v5 25.71 9.75 24.41 20.23 16.01 3
v2+v5 29.82 13.14 28.47 24.42 20.09 1
Table 2: Various README versions data performance.
In Set-2, we focus on the efficacy of different data versions when finetuning with the GPT-2 model. The results, as reflected in Table 2, indicate that high-quality expert data (v2) demonstrates clear superiority over unexamined expert data (v1), emphasizing the crucial role of ECE-examiner in enhancing data quality. Furthermore, high-quality synthetic data (v5) outperforms the unexamined expert data (v1), underscoring the significant value of data augmentation in the ECE pipeline. Notably, the combination of v2 and v5 shows improved performance over v2 alone, suggesting that including synthetic data is beneficial. This composite approach of v2+v5 leading the rank underscores the efficacy of our EAE Human-AI team.
5.3.2 Expert and Synthetic Data Integration
ROUGE1 J2L J+C2L J+G2L J+C+G2L
RANDOM 19.03 19.65 26.21 26.97
SYNTAX(R) 19.82(-2.28) 20.65(-2.78) 27.56(-8.36) 28.33(-8.84)
SEMANTIC(R) 19.74(-1.11) 19.94(-1.39) 26.42(-0.92) 27.19(-1.96)
MODEL(R) 20.18(-0.8) 20.46(-1.88) 27.9(-2.37) 28.65(-4.22)
UMLS-f J2L J+C2L J+G2L J+C+G2L
RANDOM 8.05 9.09 15.78 16.72
SYNTAX(R) 8.53(-0.11) 9.99(-1.00) 17.76(-5.11) 17.42(-5.54)
SEMANTIC(R) 8.98(-0.38) 9.07(-0.69) 15.93(-0.02) 16.74(-0.93)
MODEL(R) 9.18(-0.63) 9.4(-0.78) 18.34(-2.5) 18.45(-2.55)
Table 3: Different data selection methods performance.
In Set-3, we explored the effects of various data selection strategies on data augmentation outcomes. We found that, compared to the other two methods, the results of SEMANTIC are closer to RANDOM. Also, the SEMANTIC-based ranking remained fairly consistent between SEMANTIC and SEMANTIC_r. However, significant differences were observed in SYNTAX and SYNTAX_r, MODEL and MODEL_r. Table 3 highlights two main findings: firstly, data selection is crucial, as all methods have better performance with higher-ranked data (e.g., SYNTAX, SEMANTIC, MODEL) over lower (e.g., SYNTAX_r, SEMANTIC_r, MODEL_r). All three methods, SYNTAX, SEMANTIC, MODEL, have better results than the RANDOM baseline. Secondly, the effectiveness of a selection algorithm is proportional to the range it creates between the top and last ranking items, with a larger gap indicating a more effective selection process. This implies that with an equal volume of v5 data, the selection methods prioritize data that maximizes the benefits of augmentation and relegates the least impactful data to the lower echelon.
5.3.3 Retrieval-augmented Generation
ROUGE1 ROUGE2 ROUGEL METEOR UMLS-F1
Without data augmentation (v2)
jargon2lay 19.47 6.06 18.53 14.74 8.76
jargon+context2lay 19.40 6.40 18.38 15.12 9.24
jargon+gen2lay 26.99 10.76 25.57 21.19 16.88
jargon+context+gen2lay 27.58 11.31 26.35 21.72 17.12
With data augmentation (v2+v5 with SYNTAX)
jargon2lay 21.98 7.42 20.88 16.98 10.95
jargon+context2lay 22.13 7.90 21.04 17.54 10.71
jargon+gen2lay 29.82 13.14 28.47 24.42 20.09
jargon+context+gen2lay 29.89 13.48 28.49 24.65 20.27
Table 4: Efficacy of incorporating medical documentation context and general definition in input data.
Set-4 results underscore the significant improvement in model performance when input data is enriched with UMLS-retrieved general definitions. As illustrated in Table 4, regardless of whether we utilize only expert-annotated data or data augmentation with AI-synthetic data, including general definitions consistently enhances effectiveness. This finding confirms the value of RAG with the general definition in the lay definition generation task. Meanwhile, adding context to the input data yields a moderate impact on model performance.
5.3.4 Model Performances Against ChatGPT
Figure 5: Comparative performance analysis of DistilGPT2, BioGPT, and LLAMA2 against GPT-3.5-turbo.
In Set-5, we observed that LLAMA2-7B’s ROUGE-1 and UMLS-F1 metrics surpassed GPT-3.5-turbo in the jargon2lay task post-training. For the jargon+gen2lay setting, DistilGPT2-88M demonstrated equivalent results to GPT-3.5-turbo, while BioGPT’s performance exceeded it, and LLAMA2-7B significantly outperformed the GPT-3.5-turbo. These findings, as depicted in Figure 5, emphasize the effectiveness of open-source, mobile-adapted smaller models when appropriately fine-tuned with high-quality datasets, offering a promising avenue for deploying lightweight yet powerful NLP tools in mobile healthcare applications to help patient education.
6 Human Evaluation
6.1 Human Evaluation settings
Our human evaluation was conducted by 5 human evaluators  [^10] All of whom hold bachelor’s degrees. We randomly selected 50 pairs of (jargon, lay definitions) from the test dataset for this human evaluation. The task for evaluators was to reference the expert definitions and choose a binary preference among the following four groups of definitions: 1) DistilGPT2-J2L vs. GPT-3.5-turbo, 2) DistilGPT2-J+C+G2L vs. GPT-3.5-turbo, 3) LLAMA2-J2L vs. GPT-3.5-turbo, 4) LLAMA2-J+C+G2L vs. GPT-3.5-turbo.
6.2 Human Evaluation Results
Figure 6: Human evaluation results (win rate).
As shown in Figure 6, although the result of adding context and general definition (DistilGPT2-J+C+G2L) is better than DistilGPT2-J2L, the win rate of the two DistilGPT2 models’ output is still significantly behind the result of GPT-3.5-turbo. For LLAMA-2, generating lay definitions directly from jargon is still not as good as GPT-3.5-turbo, but adding context and general definitions is of great help. Human evaluators prefer LLAMA2-J+C+G2L more than GPT-3.5-turbo. There are some inconsistencies between the results of human evaluation and automatic evaluation. We further interviewed two medical students  [^11] Both have hospital internship experience and concluded the following conclusions to help future improvements: 1. While all our in-house systems perform satisfactorily, GPT-3.5-turbo stands out for its flexibility and user-friendliness. Particularly, it excels in elaborating complex medical terms, offering detailed explanations and practical examples for better comprehension. 2. Recent advancements Cai et al. (2023); Zhang et al. (2023b) reveal ChatGPT’s role in enhancing patient education through interactive formats like NoteAid-interactive Zhang et al. (2023a). It enables patients to actively ask questions and seek clarifications, while the AI tailors responses to aid their understanding. This interactive approach, absent in traditional dictionary-style definitions like our README dataset, calls for next-step model distillation work or further refinement in aligning the in-house system’s outputs with patient preferences. Additionally, the development of automatic metrics aligning closely with human evaluation is another critical next step.
7 Limitations and Ethical Considerations
This study provides valuable insights, but experimental results evaluated in humans demonstrate limitations of the current work and some future directions. First, better automatic evaluation metrics need to be explored to be closer to human evaluation results. Secondly, in this paper, we have only explored some heuristic data selection methods, and we need to explore more sophisticated methods in the future. In addition, the next step of the in-house system is to collect patient preferences for human alignment, which can help us generate a more user-friendly or customized lay definition. Also, we can use ChatGPT or LLAMA2-J-C-G2L to serve as the teacher and use DistilGPT2-based systems to serve as the students, performing distillation to improve the performance of the small models post current supervisor-fine-tuning on README. Finally, more interactive ways need to be considered in the future to make the in-house system more user-friendly and patient-centric.
Consider Privacy Implications, LLMs (especially third-party APIs like ChatGPT) may raise privacy concerns when conducting patient education, which may violate HIPAA regulations. In this study, we manually annotated lay definitions on publicly available MedJEx jargon terms and obtained general definitions from accessible UMLS. We also make AI-synthetic data to help training since synthetic data generation is an active field in the clinical domain especially to overcome privacy concerns Pereira et al. (2022); Shafquat et al. (2022); Mishra et al. (2023). The trained in-house system can be deployed on the patient’s mobile to avoid patient data leaving the local area, which can better protect the patient’s privacy and security. Consider Biases, LLMs trained on large amounts of text data may inadvertently capture and reproduce biases present in the data. Therefore, an in-house system trained on our data (whether expert annotation or AI synthetic) may perpetuate incorrect information or provide inaccurate answers. Finally, although we used UMLS-based RAG to reduce hallucinations, LLMs may still generate factual errors when conducting patient education.
8 Conclusions
In conclusion, our study underscores the potential of NLP to democratize medical knowledge, enabling patient-centric care by simplifying complex medical terminology. By crafting the README dataset and integrating a data-centric Human-AI collaboration, we have not only enriched the dataset quality but also broadened the horizons for training AI models in resource-constrained scenarios. Our findings reveal that even smaller, open-source models can rival the capabilities of advanced, proprietary systems like ChatGPT when fine-tuned with meticulously curated data. This research paves the way for innovative patient education tools, making strides toward a future where all patients can navigate their health information with ease and understanding.
Prompt
Examiner - 1
model = text-davinci-003(GPT - 3)
[This examiner prompt does not use context as discussed in section 3.2.1]
Decide whether the general definition is correct.
If we can generate the lay definition from the general definition then answer is yes.
term : mg
general definition : this is short for milligram which is 1/1000 of a gram usually considered a small amount.
lay definition : A tiny amount of something, usually a drug.
answer : yes
term : vitamin c
general definition : [‘A nutrient that the body needs in small amounts to function and stay healthy. Vitamin C helps fight infections, heal wounds, and keep tissues healthy. It is an antioxidant that helps prevent cell damage caused by free radicals (highly reactive chemicals). Vitamin C is found in all fruits and vegetables, especially citrus fruits, strawberries, cantaloupe, green peppers, tomatoes, broccoli, leafy greens, and potatoes. It is water-soluble (can dissolve in water) and must be taken in every day. Vitamin C is being studied in the prevention and treatment of some types of cancer.’]
lay definition : A nutrient needed by the body to form and maintain bones, blood vessels, and skin.
answer : yes
term : nodule
general definition : [‘A small lump, swelling or collection of tissue.’]
lay definition : A growth or lump that may be cancerous or not.
answer : yes
term : qd
general definition : [‘Occurring or done each day.’]
lay definition : Every day.
answer : yes
If the general definition contains many words from the term then answer is no.
term : prochlorperzine
general definition : [‘prochlorperzine’,  ’]
lay definition : A drug used to prevent or reduce nausea and vomiting.
answer : no
term : mg
general definition : [‘mg’]
lay definition : A tiny amount of something, usually a drug.
answer : no
If the lay definition can not be generated by the general definition then answer is no.
term : Virt - Vite
general definition : [‘Virt’,  - ’, The determination of the amount of Vitamin E present in a sample.’]
lay definition : A mix of vitamins. It provides vitamin B-6, vitamin B-12 and folic acid to people who do not have enough of these for good health.
answer : no
Augmenter
system_prompt = "your job is to generate a general definition of the term for a professional medical student"
model="gpt-3.5-turbo(ChatGPT API)",
messages=[
        {"role": "system", "content": system_prompt},
        {"role": "user", "content": ""},
        {"role": "user", "content": "term : incisional."},
        {"role": "assistant", "content": "general definition : An intentional cut made to an individuals body with the intent of performing a diagnostic or therapeutic intervention."},
        {"role": "user", "content": "term : PO"},
        {"role": "assistant", "content": "general definition : Of, or relating to, or affecting, or for use in the mouth.."},
        {"role": "user", "content": prompt_t}
Examiner - 2
model = text-davinci-003(GPT - 3)
exactly same as that of Examiner - 1
Table 5: All Examiner-Augmenter-Examine prompts.